# Personality Pairing Improves Human-AI Collaboration

### v1

@@ Abstract
Here we ask how AI agent “personalities” interact with human personalities, and other traits, to shape human-AI collaboration, productivity and performance. To estimate these relationships, we conducted a large-scale preregistered randomized experiment that paired 1,258 participants with AI agents that were prompted to exhibit varying levels of the Big Five personality traits. These human-AI teams produced 7,266 display ads for a real think tank, and the quality of these ads was evaluated by 1,995 independent human raters as well as in a field experiment conducted on X, which generated nearly 5 million impressions. We found, first, that personality pairing impacted teamwork quality. For example, neurotic AI improved teamwork for agreeable humans but impaired it for conscientious humans. Second, we found productivity effects of personality pairing and a “productivity-performance trade-off" in which certain pairings (e.g., agreeable human with neurotic AI) produced fewer ads but of higher quality. Third, personality pairing influenced ad quality and performance. For example, quality improved when open humans were paired with conscientious AI and when conscientious humans were paired with disagreeable AI. Some of these pairing effects were “jagged” in that they varied across text and visual tasks. For example open humans produced higher quality images but lower quality text when paired with agreeable AI. Pairing effects were also present in other human traits, like country of origin. For example, extroverted AI improved quality for Latin American workers, but degraded quality for East Asian workers. These findings demonstrate that human-AI personality alignment significantly improves collaboration, productivity, and performance and lay a foundation for future research on improving human-AI collaboration through AI personalization.
Introduction
@@
Artificial intelligence (AI) tools have gained significant attention for their ability to increase productivity and performance across a wide range of tasks. For example, chatbots powered by large language models (LLMs) reduced the time required for mid-level writing tasks by 40% while improving quality by 18% (noy2023experimental). AI-enhanced resumes increased job hiring rates by 8% (wiles2023writing). In customer support settings, AI assistance increased productivity by an average of 14% (brynjolfsson2023genai). But these benefits also exhibited considerable heterogeneity, varying across tasks (dellacqua2023navigating, ju2025collaborating), worker seniority (wang2023friend), gender (dellacqua2023navigating, humlum2024adoption), and skill levels (noy2023experimental, brynjolfsson2023genai, otis2024uneven). In clinical settings, AI assistance reduced reading time while maintaining diagnostic accuracy, but effects varied significantly across physicians and clinical contexts (chen2024impact, yu2024heterogeneity, agarwal2023combining). In financial services and customer support, human involvement in AI-based services boosted consumer confidence and material outcomes, though in other contexts AI reduced quality due to sycophancy and over-reliance (yang2025my, vaccaro2024combinations, sergeyuk2025human). Unfortunately, experiments documenting these productivity and performance gains have generally provided participants with identical AI tools, leaving open the possibility that at least some of the observed variation in performance stems from a mismatch between AI and individual human characteristics. Performance gains may therefore increase and simultaneously vary less by age, gender, or skill level, if AI is personalized for each human collaborator.
@@
Such personalization may now be possible as foundation model companies have recently begun developing, refining and deploying AI personalities to improve human-AI interaction. For example, OpenAI introduced personality customization in models like GPT-5, offering Default, Cynic, Robot, Listener, and Nerd personalities (openai2025customizing, openai2025sycophancy). Similarly, Anthropic has begun to measure undesirable personality shifts during training, such as sycophancy, a tendency for AI to overly align with user preferences (anthropic2024character, sharma2025sycophancy), or other “evil” traits, and introduced novel training methods to mitigate such personality shifts in their models. These advances enable AI agents to be tailored to users, potentially improving human-AI collaboration and ushering in a new era of AI personalization.
@@
Despite the development of AI personalities, however, no prior research has rigorously and experimentally examined their causal effects on human-AI collaboration. Such experiments are now critical to understanding whether and how tailored AI personalities may impact teamwork, productivity, and performance as the lack experimental evidence on AI personalization hinders the optimization of AI systems for collaborative tasks and undermines our understanding of the detrimental effects of mismatched AI personalities on, for example, AI psychoses.
@@
In human teams, team composition is a critical factor that promotes trust, coordination, and teamwork (zhang2023teammate, hemmer2023, woolley2023collective). Humans exhibit variation in demographic, psychological, and behavioral traits known to impact team performance (kichuk1997, peeters2006, woolley2010ci). Theories of person-team fit suggest that when team members’ personalities are compatible, collaboration becomes more effective and leads to better results (barrick1998gma, bell2007meta). As AI has evolved from a passive tool to an active collaborator (makarius2020rising, fugener2022cognitive, anthony2023, collins2024building, ju2025collaborating), tools like Cursor or Claude Code now act on behalf of and co-create alongside human users, as team members, in real time. Like humans, these AI agents also possess characteristics that influence teamwork, technology acceptance, and even psychosocial outcomes (li2025acceptance, tully2025lower, fang2025psychosocial).
@@
Unfortunately, although human and AI traits shape human-AI teamwork, their interaction remains understudied (krakowski2025human). With AI agents increasingly operating as social actors in collaborative workflows (anthony2023, gelbrich2025), it is crucial to understand whether the principles of personality alignment apply to human-AI teams and whether AI agents can be designed to exhibit personalities that complement those of their human counterparts. To address these gaps in our understanding of human-AI collaboration, we examined how AI personalities interact with human characteristics to shape teamwork, productivity, and performance.
@@
Our study leverages the unique capabilities of AI agents to vary their individual personalities (jiang2023evaluating, chen2025persona), allowing us to systematically investigate how human-AI personality alignment influences the productivity and performance of human-AI teams. Systematic prompt variation can effectively shape AI agent behavior (schulhoff2024prompt), and seminal work on OpenAI’s ChatGPT architecture demonstrates that prompting enables control of agent behavior without task-specific fine-tuning (brown2020language). Advanced prompting methods, such as chain-of-thought prompting, have further demonstrated significant improvements in task performance by guiding a model’s reasoning processes (wei2023chainofthought) and “persona vectors" have been shown to successfully monitor and control personality shifts in AI agents (chen2025persona). These methods enable us to intentionally design AI agents with particular social and behavioral characteristics, which can be tailored to complement human collaborators. Randomized controlled trials (RCTs) are now needed to evaluate the effects of such AI personalities on human-AI collaboration and task outcomes to understand how AI agents can be personalized for effective teamwork (jakesh2023opinionated).
@@
To investigate how personality alignment between humans and AI agents influences collaboration, we conducted a large-scale experiment involving 1,258 participants representative of the U.S. population using Pairit, an experimental platform we developed for extensible human-AI collaboration (ju2025collaborating). After taking a 10-item pre-task survey to measure their Big Five personality traits, participants, who were recruited via Prolific from October 15 to 18, 2024, were randomly paired with multi-modal AI agents, which had full context of the workspace and were induced to independently exhibit high or low levels of the Big Five personality traits using P2 prompting (jiang2023evaluating). These human-AI teams then worked together to create display ads marketing the year-end report of a real think tank within a 40-minute session, using real-time chat and synchronized text and image editing tools. Teams were provided with image assets by the think tank and could separately call Dall-E 3 to generate new images for the campaigns. As described in the Materials and Methods, the AI agents were fully multi-modal, had access to the Internet, could manipulate the images, the image carousel and the text of the ads, text with their human collaborators and could also call Dall-E 3 to generate new images. This setup allowed us to assess the causal effects of personality alignment on collaboration outcomes. The randomization of AI personalities, such as high or low conscientiousness, openness or extroversion, enabled us isolate the causal effects of human-AI personality pairing on the collaboration, productivity and performance of human-AI teams.
@@
Importantly, we approach personality prompts as neither about memory, such as recalling past chat interactions, nor about aligning LLM responses to human preferences which often leads to AI sycophancy (chhikara2025mem0, sharma2025sycophancy). Instead, we randomize AI personalities to learn the causal effects of various personality pairings on human-AI team performance. This approach parallels findings in human-robot interaction, where robots’ social communication and appearance have been shown to influence team performance (jung2013backchanneling, jung2018theories).
To assess the quality of the display ads produced by the human-AI teams, we conducted a separate evaluation study using a second independent sample of 1,198 new participants recruited from Prolific. These raters were based in the US and selected to ensure representation across gender, age, and ethnicity. Each rater evaluated a random sample of 40 ads, with the sampling design guaranteeing that every ad received at least three independent ratings. Raters assessed each ad on three dimensions using 7-point Likert scales: text quality (“The text is present, clear, relevant, and engaging”), image quality (“The image is visually appealing”), and likelihood of clicking the ad (“I am likely to click on this ad”). The ads were presented as realistic mockups, incorporating standard display ad elements such as images, copy, call-to-action buttons, and social media-style interaction options (e.g., Like, Comment, Share; see Supporting Information [SI] Section C).
To evaluate the real-world performance of the display ads, we then conducted a pre-registered field experiment on the X platform. We sampled 2,000 ads from the total 7,266 ads using stratified sampling based on predicted click likelihoods implied by the human ratings to ensure sampling across the ad quality distribution. These ads were then run in 400 geographically non-overlapping campaigns, each with a five-ad split test (braun2024abtesting, braun2024adplatforms). Key performance metrics included the click-through rate (CTR) and the cost-per-click (CPC) measured on X, as well as the view-through rate (VTR; fraction of document viewed) and the view-through duration (VTD; seconds) tracked via unique DocSend links for each ad (details of the field experiment are provided in the Materials and Methods).
@@ Results
Our findings revealed that personality pairings significantly affected collaborative outcomes including teamwork quality, productivity, ad quality and ad performance. For example, humans high in conscientiousness created ad text of greater quality when paired with conscientious AI agents, while extroverted humans paired with conscientious AI produced lower quality text, lower quality images, and ads with lower click likelihood. Some pairings demonstrated uneven “jagged” performance patterns, with effects varying sharply between the textual and visual components of the task. We further observed effects on productivity, with agreeable humans, for example, producing more ads when paired with open AI. Additionally, certain combinations yielded a performance-productivity trade-off, such as agreeable humans with neurotic AI generating fewer ads, but ads of higher quality. Overall, these results underscore the substantial role of personality alignment in human-AI teamwork and lay a foundation for future research on AI personalization to improve human-AI productivity and collaborative performance. In particular, we found specific effects on teamwork, productivity and output quality.
Teamwork
First, we found personality pairings significantly influenced teamwork quality. To evaluate these effects, participants completed a 35-item teamwork quality survey after the task, which assessed six facets of collaboration: communication, coordination, balance of member contributions, mutual support, effort, and cohesion (hoegl2001). We conducted a regression analysis on data produced by our experiments to identify the average treatment effects of AI personalities as well as the interaction effects of human and AI personality traits in combination and found statistically significant effects for human openness, conscientiousness, and agreeableness, as well as for AI neuroticism (Figure 2, see SI Section H).
Figure 2: Regression coefficients for human-AI personality interaction terms for teamwork quality (p<<0.05) (hoegl2001). Error bars show 95% confidence intervals. See SI Section H for the full regression tables.
In examining interaction effects, we found that for open humans, conscientious AI lowered the balance of member contributions, mutual support, effort, and cohesion. For conscientious humans, open AI improved balance and cohesion, and conscientious AI improved effort and cohesion. However, for conscientious humans, neurotic AI lowered communication, coordination, balance, and mutual support. Finally, for agreeable humans, neurotic AI improved communication, coordination, mutual support, and effort. However, these positive results reversed when agreeable humans were paired with open AI (for cohesion) and with conscientious AI (for coordination, balance, and cohesion). These effects revealed that personality pairings shape perceived teamwork dynamics in human-AI collaboration.
Ad quality
In human evaluations of ad quality, we found that personality pairing between humans and AI agents significantly influenced ad text quality (Figure 3) beyond the effects of AI personality alone (see SI Section H). The average treatment effects demonstrated that conscientious AI reduced, and extroverted AI improved, text quality on average. The interaction effects then demonstrated statistically and economically significant effects of human-AI personality pairing on ad quality. For both open and conscientious humans, conscientious AI increased text quality, but agreeable AI decreased text quality. For extroverted humans, extroverted AI increased text quality, but conscientious AI decreased text quality. For agreeable humans, neurotic AI increased text quality, but extroverted AI decreased text quality. These results indicate that aligning AI personality traits with human traits impacts the effectiveness of collaborative ad text production.
Figure 3: Regression coefficients for human-AI personality interaction terms on ad text quality, image quality, and click likelihood (p<<0.05). Error bars show 95% confidence intervals. See SI Section H for the full regression table.
For image quality, we observed a different set of significant interaction effects based on human-AI personality pairing. For open humans, agreeable AI increased image quality while, for conscientious humans, open AI increased image quality. For extroverts and agreeable participants, being paired with conscientious and open AI, respectively, decreased image quality.
For click likelihood ratings, which reflect the perceived engagement potential of the ads, we observed no significant effects for humans with high openness or agreeableness. For conscientious humans, agreeable AI decreased click likelihood, while for extroverted and neurotic humans, conscientious AI decreased click likelihood.
Across the three measures of ad quality (text quality, image quality and perceived click likelihood), personality pairings between humans and AI agents consistently influenced outcomes. For extroverts, conscientious AI decreased quality across all measures. For conscientious humans, agreeable AI decreased text quality and click likelihood. Some pairings had significant and opposite effects, indicating a “jagged edge” of human-AI competence across tasks and personality pairings (dellacqua2023navigating). For example, for open humans, agreeable AI improved image quality but worsened text quality. Collectively, these patterns suggest that strategic personality pairings can optimize human-AI collaboration outcomes, with some generalizable and some task-specific personality pairings.
In addition to participant personality traits, we found that participant demographics, including previous AI experience, employment status, and country of origin, moderated the effects of AI personality traits on ad quality. For individuals with previous AI experience, extroverted AI hurt overall performance, while neurotic AI improved performance (see SI Sections G).
Although all participants were based in and representative of the demographic composition of the U.S., their diverse countries of birth allowed us to explore possible cultural and regional differences in collaborative outcomes (Figure 4, see SI Section H). Participants from Western regions, the Middle East & South Asia produced significantly lower quality images when paired with agreeable AI, while participants from East Asia produced images of significantly lower quality when paired with extroverted and agreeable AI. Participants from Latin America produced significantly higher quality ad text when paired with extroverted AI and significantly lower quality ad text when paired with conscientious, agreeable, and neurotic AI. For participants from all regions except East Asia, agreeable AI significantly lowered click likelihood. For those from Latin America, extroverted AI improved click likelihood, while conscientious AI lowered click likelihood. For those from the Middle East & South Asia, agreeable AI lowered click likelihood.
Interestingly, we observed that agreeable AI consistently undermined performance across all measures, reducing image quality (for Western, East Asian, Middle Eastern & South Asian workers), text quality (for Latin American workers), and click likelihood (for Western, Latin American, Middle Eastern & South Asian workers). Taken together, these patterns suggest individual characteristics like culture mediate how AI personality traits, especially agreeable or sycophantic AI (sharma2025sycophancy), affect collaboration and suggest the potential for future research investigating the strategic pairing of AI personalities with various human traits.
Figure 4: Regression coefficients for interaction terms of a participants’ country of birth and AI personality on ad quality measures (p<<0.05). Error bars show 95% confidence intervals. See SI Section H for the full regression table.

### v2

@@ Abstract
Here we examine how AI agent “personalities” interact with human personalities to shape human-AI collaboration and performance. In a large-scale, preregistered randomized experiment, we paired 1,258 participants with AI agents prompted to exhibit varying levels of the Big Five personality traits. These human-AI teams produced 7,266 display ads for a real think tank, which we evaluated using 1,995 independent human raters and a field experiment on X that generated nearly 5 million impressions. We found that human and AI personalities individually shaped ad quality and teamwork. When examined together, human-AI personality pairings directly effected ad quality outcomes. For example, extraverted humans paired with conscientious AI produced the lowest-quality ads, followed by conscientious humans paired with agreeable AI and neurotic humans paired with conscientious AI. In the field experiment, ad quality significantly influenced ad performance, measured by click-through rates and cost-per-click, and neurotic humans paired with neurotic AI achieved higher click-through rates, even after controlling for ad quality. Together, these results provide the first large-scale causal experimental evidence that specific personality pairings can improve human-AI collaboration and motivate future research on the implications of AI personalization for performance and teamwork dynamics in human-AI teams.
@@
Artificial Intelligence (AI) tools have gained significant attention for their ability to increase productivity and performance across a wide range of tasks. For example, chatbots powered by large language models (LLMs) reduced the time required for mid-level writing tasks by 40% while improving quality by 18% (Noy and Zhang, 2023). AI-enhanced resumes increased job hiring rates by 8% (Wiles et al., 2023). In customer support settings, AI assistance increased productivity by an average of 14% (Brynjolfsson et al., 2023). But these benefits also exhibited considerable heterogeneity, varying across tasks (Dell’Acqua et al., 2023, Ju and Aral, 2025), worker seniority (Wang et al., 2023), gender (Dell’Acqua et al., 2023, Humlum and Vestergaard, 2024), and skill levels (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Otis et al., 2024). In clinical settings, AI assistance reduced reading time while maintaining diagnostic accuracy, but effects varied significantly across physicians and clinical contexts (Chen et al., 2024, Yu et al., 2024, Agarwal et al., 2023). In financial services and customer support, human involvement in AI-based services boosted consumer confidence and material outcomes, though in other contexts AI reduced quality due to sycophancy and over-reliance (Yang et al., 2025, Vaccaro et al., 2024, Sergeyuk et al., 2025). Unfortunately, experiments documenting these productivity and performance gains have generally provided participants with identical AI tools, leaving open the possibility that at least some of the observed variation in performance stems from a mismatch between AI and individual human characteristics. Performance gains may therefore increase and simultaneously vary less by age, gender, or skill level, if AI is personalized for each human collaborator.
@@
Such personalization may now be possible as foundation model companies have recently begun developing, refining and deploying AI personalities to improve human-AI interaction. For example, OpenAI introduced personality customization in models like GPT-5, offering Default, Cynic, Robot, Listener, and Nerd personalities OpenAI (2025b, a). Similarly, Anthropic has begun to measure undesirable personality shifts during training, such as sycophancy, a tendency for AI to overly align with user preferences Anthropic (2024), Sharma et al. (2025), and introduced novel training methods to mitigate such personality shifts in their models. These advances enable AI agents to be tailored to users, potentially improving human-AI collaboration and ushering in a new era of AI personalization.
@@
Despite the development of AI personalities, however, no prior research has rigorously and experimentally examined their causal effects on human-AI collaboration. Such experiments are now critical to understanding whether and how tailored AI personalities may impact teamwork and performance as the lack of experimental evidence on AI personalization hinders the optimization of AI systems for collaborative tasks and undermines our understanding of the detrimental effects of mismatched AI personalities on, for example, adverse psychosocial outcomes (Fang et al., 2025).
@@
In human teams, team composition is a critical factor that promotes trust, coordination, and teamwork (Zhang et al., 2023, Hemmer et al., 2023, Woolley et al., 2023). People exhibit variation in demographic, psychological, and behavioral traits known to impact team performance (Kichuk and Wiesner, 1997, Peeters et al., 2006, Woolley et al., 2010). Theories of person-team fit suggest that when team members’ personalities are compatible, collaboration becomes more effective and leads to better results (Barrick et al., 1998, Bell, 2007). As AI has evolved from a passive tool to an active collaborator (Makarius et al., 2020, Fügener et al., 2022, Anthony et al., 2023, Collins et al., 2024, Ju and Aral, 2025), tools like Cursor or Claude Code now act on behalf of and co-create alongside human users, as team members, in real time. Like humans, these AI agents also possess characteristics that influence teamwork, technology acceptance, and even psychosocial outcomes (Li et al., 2025, Tully et al., 2025, Fang et al., 2025).
@@
Early work on computers as social actors showed that people prefer personality-matched computers (Nass and Moon, 2000), and personality-targeted advertising can alter behavior at scale (Matz et al., 2017, 2024). However, these studies examined either preference outcomes with pre-LLM systems or one-shot content personalization, not interactive collaboration with AI agents. The interaction between human and AI traits in collaborative production remains understudied Krakowski et al. (2025). With AI agents increasingly operating as social actors in collaborative workflows (Anthony et al., 2023, Gelbrich et al., 2025), it is crucial to understand whether the principles of personality pairing apply to human-AI teams and whether AI agents can be designed to exhibit personalities that complement those of their human counterparts. To address these gaps in our understanding of human-AI collaboration, we examined how AI personalities interact with human characteristics to shape teamwork and performance.
@@
Our study leverages the unique capabilities of AI agents to vary their individual behavioral styles (Jiang et al., 2023, Chen et al., 2025), allowing us to investigate how human-AI personality pairing influences the performance of human-AI teams. Systematic prompt variation can effectively shape AI agent behavior (Schulhoff et al., 2024), and seminal work on OpenAI’s ChatGPT architecture demonstrates that prompting enables control of agent behavior without task-specific fine-tuning (Brown et al., 2020). Advanced prompting methods, such as chain-of-thought prompting, have further demonstrated significant improvements in task performance by guiding a model’s reasoning processes (Wei et al., 2022) and “persona vectors" have been shown to successfully monitor and control personality shifts in AI agents (Chen et al., 2025). These methods enable us to intentionally design AI agents with particular social and behavioral characteristics, which can be tailored to complement human collaborators. Randomized controlled trials (RCTs) are now needed to evaluate the effects of such AI personalities on human-AI collaboration and task outcomes to understand how AI agents can be personalized for effective teamwork (Jakesch et al., 2023).
@@
To investigate how personality pairing between humans and AI agents influences collaboration, we conducted a large-scale experiment involving 1,258 participants representative of the U.S. population using Pairit, an experimental platform we developed for extensible human-AI collaboration (Ju and Aral, 2025). After taking a 10-item pre-task survey to measure their Big Five personality traits, participants, who were recruited via Prolific from October 15 to 18, 2024, were randomly paired with multi-modal AI agents, which had full context of the workspace and were induced to independently exhibit high or low levels of the Big Five personality traits using P2 prompting (Jiang et al., 2023). These human-AI teams then worked together to create display ads marketing the year-end report of a real think tank during a 40-minute session, using real-time chat and synchronized text and image editing tools. Teams were provided with image assets by the think tank and could separately call Dall-E 3 to generate new images for the campaigns.
Throughout the collaboration, the AI agent dynamically received screenshots of the current ad image alongside the task brief, the current ad copy, the chat history with the human, and its own action history. As it was provided this context, it could choose from five action types: Wait, Chat, EditText (revise the headline, primary text, description, or image prompt), SelectImage (choose an image from the carousel), or GenerateImage (call Dall-E 3 to produce a new image). The agent and the human shared a single live workspace, so the human could see and edit the same ad in real time as the agent acted. Only the human partner could submit ads, and teams iterated to submit as many ads as they could within the 40-minute window. This setup allowed us to assess the causal effects of personality pairing on collaboration outcomes. The randomization of AI agent personalities, such as high or low conscientiousness, openness or extroversion, enabled us isolate the causal effects of human-AI personality pairing on the collaboration and performance of human-AI teams.
@@
Importantly, we approached personality prompts as neither about memory, such as recalling past chat interactions, nor about aligning LLM responses to human preferences which often leads to AI sycophancy (Chhikara et al., 2025, Sharma et al., 2025). Instead, we treated AI personality as an experimental treatment: each AI agent was independently and randomly assigned a personality profile (high or low on each Big Five personality trait), and we estimated how this assignment interacted with the human partner’s measured traits to shape collaboration outcomes. Random assignment rules out selection effects, such as users gravitating toward AI styles they prefer, so observed differences across pairings reflect causal effects of the pairing itself. This approach parallels findings in human-robot interaction, where robots’ social communication and appearance have been shown to influence team performance (Jung et al., 2013, Jung and Hinds, 2018).
To assess the quality of the display ads produced by the human-AI teams, we conducted a separate evaluation study using a second independent sample of 1,198 new participants recruited from Prolific. These raters were based in the US and selected to ensure representation across gender, age, and ethnicity. Each rater evaluated a random sample of 40 ads, with the sampling design guaranteeing that every ad received at least three independent ratings. Raters assessed each ad on three dimensions using 7-point Likert scales: text quality (“The text is present, clear, relevant, and engaging”), image quality (“The image is visually appealing”), and likelihood of clicking the ad (“I am likely to click on this ad”). The ads were presented as realistic mockups, incorporating standard display ad elements such as images, copy, call-to-action buttons, and social media-style interaction options (e.g., Like, Comment, Share; see Materials and Methods).
To evaluate the real-world performance of the display ads, we then conducted a pre-registered field experiment on the X platform. We sampled 2,000 ads from the 7,266 total ads produced during the experiment, using stratified random sampling based on the predicted click likelihoods implied by the human ratings to ensure sampling across the ad quality distribution. These ads were then run in 400 geographically non-overlapping campaigns, each with a five-ad split test (Braun and Schwartz, 2024, Braun et al., 2024). Key performance metrics included the click-through rate (CTR) and the cost-per-click (CPC) measured on X, as well as the view-through rate (VTR; fraction of document viewed) and the view-through duration (VTD; seconds) tracked via unique DocSend links for each ad (details of the field experiment are provided in the Materials and Methods). We employed ZIP-code-level randomization across campaigns to prevent audience spillovers; campaign random effects were specified to absorb campaign-level variation in audience composition and temporal confounders; and controls for advertising spend accounted for any remaining differences in budget allocations across campaigns and ads created by auto-bidding. We applied Benjamini-Hochberg (BH) multiple hypothesis testing corrections to all analyses in the paper and report results based on these corrections throughout.
@@ Results
Our results demonstrate that personality pairing affected ad quality, teamwork, and real-world ad performance in our experiments. We first report the individual effects of human and AI personality traits on teamwork and ad quality, then turn to the personality pairing results and conclude by reporting the results of the field experiment.
Individual trait effects on teamwork and ad quality
We first tested whether measured human traits and randomized AI traits independently predicted teamwork and ad quality in human-AI teams (Figure 2). Each model regressed an outcome on the five human trait scores and the five AI trait scores with no interaction terms, and BH corrections were applied within each outcome family. Since AI traits were experimentally randomized, while human traits were measured but not manipulated, the AI coefficients can be interpreted causally, whereas the human trait coefficients represent correlations not causal relationships.
Figure 2: Effect of personality on ad quality and teamwork. Coefficients from regressions of each outcome on the five human trait scores and the five AI trait scores, with no interaction terms, BH-corrected within each outcome family. Panel A shows ad quality dimensions (text, image, click likelihood); panel B shows teamwork indices (task, social). Within each panel, human trait effects appear on the left and AI trait effects on the right. X-axis labels use O = Openness, C = Conscientiousness, E = Extraversion, A = Agreeableness, and N = Neuroticism. Symbols indicate significance after BH correction: $^{\dagger}p<0.10$, $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$. Error bars show 95% confidence intervals. See SI Appendix, Section LABEL:si:full, for the full regression tables.
Human traits predicted both teamwork and ad quality. Conscientious and agreeable humans experienced stronger task ($p<0.001$, $p<0.01$) and social teamwork ($p<0.001$, $p<0.001$), while neurotic humans reported weaker task teamwork ($p<0.05$). Conscientiousness and agreeableness in humans also predicted higher ratings on all six underlying teamwork dimensions (communication, coordination, balance of contributions, mutual support, effort, and cohesion) (see SI Appendix, Section LABEL:si:main-effects). Human openness predicted higher text quality ($p<0.05$) and human extraversion predicted lower text quality ($p<0.05$). AI traits also affected ad quality: AI extraversion increased text quality ($p<0.05$), whereas AI conscientiousness reduced text quality ($p<0.01$) (see SI Appendix, Section LABEL:si:main-effects, for the full regression tables). It is perhaps not surprising that human and AI personalities affect teamwork and ad quality. But evaluating the potential benefits of AI personalization requires a more comprehensive analysis of human-AI personality pairings.
Figure 3: Effect of personality pairing on ad quality. Coefficients for human-AI personality interaction terms on the composite ad quality index (gray) and the three ad quality dimensions: text quality (blue), image quality (red), and click likelihood (green), BH-corrected within each outcome family. Panel titles show the full human trait names; x-axis labels use O = Openness, C = Conscientiousness, E = Extraversion, A = Agreeableness, and N = Neuroticism. Symbols indicate significance after BH correction: $^{\dagger}p<0.10$, $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$. Error bars show 95% confidence intervals. See SI Appendix, Section LABEL:si:full, for the full regression table.
Personality pairing and ad quality
Beyond individual traits, personality pairings between humans and AI agents significantly influenced ad quality (preregistered; Figure 3; see SI Appendix, Section LABEL:si:prereg). Three pairings drove the strongest effects on the composite ad quality index and all three reduced quality. Extraverted humans paired with conscientious AI produced the lowest-quality ads ($p<0.001$), followed by conscientious humans paired with agreeable AI ($p=0.01$) and neurotic humans paired with conscientious AI ($p=0.049$).
These results suggest a broad pattern in personality pairing effects on human-AI collaboration: when humans who are assertive, goal driven and set in their ways (traits associated with extraversion and conscientiousness) collaborate with compliant AI that are not skeptical of and do not push back on human ideas, output quality decreases. For example, when extraverted humans, who are assertive, enthusiastic and “high energy," collaborate with conscientious AI, the result is a collaboration that efficiently and effectively implements the human’s plans, with little explicit push back from AI on whether they are the right plans. When conscientious humans, who are disciplined and goal-directed, collaborate with agreeable AI, the result is a more compliant implementation of human ideas. When neurotic humans, who are more reactive to stress and uncertainty, collaborate with conscientious AI, the team implements a more volatile and less stable collaborative process, again with limited explicit push back from the AI, resulting in lower quality output.
Exploratory LLM labeling of the human-AI chat logs is consistent with this interpretation. Relative to the corresponding low-trait AI comparison cells, conscientious AI paired with extraverted humans showed slightly less AI questioning (0.331 vs. 0.335) and slightly more supportiveness (0.382 vs. 0.376), while the humans themselves were slightly more assertive (0.213 vs. 0.202). Agreeable AI paired with conscientious humans showed lower AI push back (0.003 vs. 0.004) and higher order-taking (0.107 vs. 0.104), while the humans themselves were more goal-oriented (0.239 vs. 0.212). Conscientious AI paired with neurotic humans again showed lower AI push back (0.004 vs. 0.005) and higher order-taking (0.114 vs. 0.109) (see SI Appendix, Section LABEL:si:sycophancy). Overall, explicit AI push back was rare in the human-AI chats ($1\%$ of messages). This suggests that recent trends toward AI sycophancy are likely to produce lower quality outcomes in human-AI collaboration and that extraverted, conscientious and neurotic humans are likely to be the most vulnerable to performance degradation when collaborating with sycophantic AI. Although this linguistic analysis is exploratory and not necessary to confirm the conclusions of our statistical analysis, that linguistic indicators conform to the patterns in our statistical results increases our confidence in our interpretations and conclusions.
At the level individual performance indicators, extraverted humans paired with conscientious AI lowered quality across all three measured dimensions including text ($p<.001$), image ($p<.001$), and click likelihood ($p<.05$), while pairing with extraverted AI increased text quality ($p<.01$). Conscientious humans produced lower quality text when paired with agreeable AI ($p<.001$) and higher quality text and images when paired with conscientious ($p<.01$) and open AI ($p<.05$). Open humans produced higher quality text when paired with conscientious AI ($p<.05$) and lower quality text when paired with agreeable AI ($p<.05$). Agreeable humans produced lower quality text when paired with extraverted AI ($p<.05$) and lower quality images when paired with open AI ($p<.05$). Finally, neurotic humans produced ads with lower click likelihood when paired with conscientious AI ($p<.05$).
AI experience (preregistered) moderated how AI personality affected ad quality, while employment status and country of origin showed additional exploratory heterogeneity (see SI Appendix, Section LABEL:si:demographics; see also SI Appendix, Section LABEL:si:prereg). Taken together, these results suggest the existence of economically and statistically significant gains from AI personalization. When people collaborate with AI with personalities that are more appropriate for them, the quality of their output improves significantly.
Personality pairing and teamwork
Personality pairings also shaped the social side of teamwork, though more weakly than ad quality (preregistered; see SI Appendix, Section LABEL:si:prereg). Participants completed a 35-item teamwork quality survey after the task (Hoegl and Gemuenden, 2001), which assessed six facets of teamwork grouped into a social dimension (mutual support, effort, and cohesion) and a task dimension (communication, coordination, and balance of member contributions). The social dimension showed joint significance of personality pairing effects ($F$-test $p=0.022$; MANOVA $p=0.004$), although no individual interactions survived the BH corrections, and the task dimension did not show joint significance. These results suggest that personality pairings may influence collaboration dynamics, though individual effects require confirmation in future targeted experiments. See SI Appendix, Section LABEL:si:full, for the full coefficient plots and regression tables.